OpenStudy (anonymous):

Consider the function f(x)=x^(2)e^(x)-2xe^(x)-3e^(x). (A) Find the equation of the normal line to f where x=0. (B) Find the value(s) of x for which f has a horizontal tangent.

OpenStudy (turingtest):

can you find f' ?

OpenStudy (anonymous):

Yeah would it be 2xe^(x)-2e^(x)-e^(x)

OpenStudy (anonymous):

f(x)=x^(2)e^(x)-2xe^(x)-3e^(x). f'(x)=e^x(x^2-5) for horizontal tangents f'(x)=e^x(x^2-5)=0 ---> x1=sqrt5=2.236 x2=-sqrt5=-2.236 plug them to orig eq to find your y

OpenStudy (anonymous):

your normal line would be perpendicular to these tangent line

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

When plugging in the x-values what do I do with the e's.

OpenStudy (anonymous):

plug in also the x values to e

OpenStudy (anonymous):

Do I do both x-values at once or separately?

OpenStudy (anonymous):

use your calculator

OpenStudy (anonymous):

f(x)=sqrt5^(2)e^(sqrt5)-2sqrt5e^(sqrt5)-3e^(sqrt5) this is also f(x)=e^x (x^2-2x-3) plug in the x here

OpenStudy (anonymous):

f(x)=e^x (x^2-2x-3) f(sqrt5)=e^sqrt5 ((sqrt5)^2-2sqrt5-3)=-23 f(sqrt-5)=e^sqrt-5 ((sqrt-5)^2-2sqrt-5-3)=0.692 P(x,y)=P(sqrt5,-23)., Q(x,y)=Q(-sqrt5,0.692) these are your horizontal tangents

OpenStudy (anonymous):

next is to find the normal line of f(x) at x=0

OpenStudy (anonymous):

Yay I got the same answers as you :)

OpenStudy (anonymous):

ok did you find the normal line eq?

OpenStudy (anonymous):

very good

OpenStudy (anonymous):

No what do I do to get that?

OpenStudy (anonymous):

the normal line to f(x) will have a perpendicular to that line meaning the slope has a formula m=-1/m (slope for perpendicular line)

OpenStudy (anonymous):

ist get your alope at x=0

OpenStudy (anonymous):

slope m=f'(x)=e^x(x^2-5) now at x=0 m=e^0(0^2-5)=-5 now the line perpendicular to it will have a slope of m=-1/m m=-1/-5=1/5 also you y value at x=0 is y=f(0)=e^0 (0-2(0)-3) y=-3 now you have (0,-3) your equation normal to the line is y-y1=m(x-x1) y-(-3)=1/5 (x-0) y+3=1/5 x y=1/5 x -3 Normal line to f(x) ans

OpenStudy (anonymous):

did you get the idea? any question,

OpenStudy (anonymous):

Yeah thanks you helped immensely.Now I can do the rest of my homework problems.

OpenStudy (anonymous):

very good,,good luck now..have fun ..lol :D

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw